OpenStudy (anonymous):

What is the derivative of h(a) = (4acosa)^2

OpenStudy (anonymous):

I keep getting h'(a) = 8(cosa - asina) but the answer (according to WolframAlpha) is y'=-32(x^2 sinxcosx - xcos(^2)x)

hartnn (hartnn):

(16 a^2 cos^2 a)' =16 [2a cos^2 a - a^2 (2cos a sin a) ] =......

OpenStudy (anonymous):

@hartnn I got = 32(acos^2a - a^2sinacosa) in the end. I forgot to use the g'(x) portion of the chain rule in this problem! Silly me. But thank you.